OpenStudy (anonymous):

How would you sketch a graph with the y-intercept being -3, the x-intercept being 1 and -3, and i think the vertex is1.

OpenStudy (lgbasallote):

it cannot be plotted...considering as your vertex isnt even a coordinate lol

OpenStudy (anonymous):

how would you find the vertex?

OpenStudy (lgbasallote):

uhmm idk..you just gave it lol

OpenStudy (anonymous):

i thought that was it im not sure - the problem is y=x^2+2x-3 so I am not sure

OpenStudy (lgbasallote):

hmm i am assuming this is a parabola...

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

ahhh...hint..rewrtie it in the form \(y = a(x+h)^2 + k\) where \((h, k) = \text{vertex}\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

 the vertex is (-1,-4) the graph is y+4=(x+1)^2